C=160+0.6Yd

I=150

G=150

T=100

Yd=Y-T

(a) Assume that output is equal to 900. Compute total demand. Is it equal to output? Explain.

(b) Assume that output is equal to 1000. Compute total demand. Is it equal to output? Explain.

(c) Assume output is equal to 1000. Compute private saving. Is it equal to investment? Explain.
Answer #1

a. Total Demand is the sum of C+I+G

= 160+0.6*(900-100)+150+150 = 940

No, it is more than output and the reason is because of inefficient production levels

b. 160+0.6*(1000-100)+150+150 = 1000.

Yes. As the economy is producing at its optimal levels

c. private saving = Y-C-T = 1000-(160+0.6*(1000-100))-100 =1000-700-100 =200

No it is greater than Investment as the people have less confidence in the economy and would save than investing